OpenStudy (anonymous):

Chain Rule Calculus f(x)=(2x-7)^3 I NEEED HELP!!

OpenStudy (bahrom7893):

3(2x-7)^2 * (2*1-0)

OpenStudy (anonymous):

may i ask how you got that?

OpenStudy (bahrom7893):

(f(x))^a = a(f(x))^(a-1) * f'(x)

OpenStudy (anonymous):

f'(x) = 3(2x-7)^2 * d/(dx (2x-7) =3(2x-7)^2(2) =6(2x-7)^2

OpenStudy (anonymous):

cant u jst simplify tis using formula for (a-b)^3=a^3+3ab^2-3a^2b-b^3 and then diff tis.. its much easier than to use product rule